# On the Crucial Role of Initialization for Matrix Factorization

### v1

@@ 1 Introduction
ScaledGD (Jia et al., 2023) small ${\cal O}(\log(1/\epsilon))$
@@
OP AltGD (Ward and Kolda, 2023) special ${\cal O}\big{(}\kappa^{2}\log(1/\epsilon)\big{)}$
@@ 1.1 Related works
We only streamline results on the convergence of matrix factorization under the broad umbrella of optimization under forth order growth, where the objective is to minimize a forth-order polynomial with first order approahces. Other closely related topics, such as LoRA variants, can be found in Apdx. A.1.
Optimization under forth-order growth. Matrix factorization problems considered in this work are classical examples of forth-order growth functions. It involves a complex landscape characterized by nonconvexity, nonsmoothness, and the absence of PL condition. Similar to other works listed in Tab. 1, the goal of this work is to recap this classical problem and to unveil intriguing behaviors from an optimization perspective. Recent works have examined the convergence of several algorithms, such as GD, AltGD, and ScaledGD (Du et al., 2018; Ye and Du, 2021; Ward and Kolda, 2023; Jia et al., 2023; Jiang et al., 2023b) in exact- and over- parametrized settings. Most of them admit linear convergence with different dependences on the condition number of the factorized matrix $\mathbf{A}$. GD for matrix square root problems is studied in (Jain et al., 2017). Another closely related setting within forth-order growth is matrix sensing; see e.g., (Stöger and Soltanolkotabi, 2021; Zhang et al., 2021; Jin et al., 2023; Xiong et al., 2024; Xu et al., 2023). Linear rates are obtained for problems with exact- and over- parametrization, despite some of them demand early stopping. Similar to matrix factorization, not too much is known for under-parametrization. There are other approaches to tackle general forth-order growth optimization. For example, relative smoothness is considered in (Lu et al., 2018); adaptive step sizes induced by fine-grained geometry are studied in (Davis et al., 2024). The work of (Dragomir and Nesterov, 2023) also copes with such problems but requires convexity of the objective.
@@ 2.1 Preliminaries
Within this section, we assume that $\mathbf{A}\in\mathbb{R}^{m\times m}$ is positive semidefinite (PSD), otherwise one can employ the asymmetric formulation as in later sections. Problem (1) also closely links with matrix sensing, particularly under a sufficient number of Gaussian measurements (Xiong et al., 2024). From an optimization perspective, problem (1) is nonconvex and has no global Lipschitz gradient. These undesirable properties pose challenges for analyzing the convergence of classical optimization approaches (Tu et al., 2016; Chi et al., 2019).
@@ 2.2 Nyström initialization
where $\mathbf{\Omega}\in\mathbb{R}^{m\times r}$ is a Gaussian random matrix. From this initialization, it is not difficult to see that condition i) is satisfied already. Our next lemma shows that the condition ii) holds w.h.p.
@@
The detailed expression for this “high probability” in Lemma 1 can be found in Apdx. B.1.1. Note that there will be a “w.h.p.” over the initialization in most of our results. This refers to that $\text{rank}(\mathbf{X}_{0})=r$ in needed for exact- and under-parametrized settings, and $\text{rank}(\mathbf{X}_{0})=r_{A}$ is needed when over-parametrized.
@@ 3.2 Nyström initialization in the exact-parametrized setting
With Nyström initialization (5) and different choices of step sizes, modified ScaledGD in (6) have a similar behavior as Theorem 1, that is, a two-phase behavior w.h.p. over the initialization:
@@ 4 NoRA: Nyström low rank adapters
where $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$ is the pretrained weight (of a particular layer), and $\mathbf{X}\in\mathbb{R}^{m\times r}$ and $\mathbf{Y}\in\mathbb{R}^{n\times r}$ with $r\ll\min\{m,n\}$. A more detailed recap of LoRA can be found in Apdx. A.1. Directional alignment can be achieved if singular vectors for $\Delta\mathbf{W}$ are leveraged to initialize $\mathbf{X}_{0}$ and $\mathbf{Y}_{0}$. While $\Delta\mathbf{W}$ is unavailable a priori, empirical wisdom suggests that there exists a set of well-performed adapters that lie in the column (row) span of the pretrained weight matrix (Lingam et al., 2024), i.e., $\text{ColSpan}(\Delta\mathbf{W})\subseteq\text{ColSpan}(\mathbf{W}_{0})$ and $\text{RowSpan}(\Delta\mathbf{W})\subseteq\text{RowSpan}(\mathbf{W}_{0})$. In other words, $\mathbf{W}_{0}$ can be adopted as a suitable replacement of $\Delta\mathbf{W}$ for directional alignment.
@@
We note that ScaledGD has already been applied for LoRA training in (Zhang and Pilanci, 2024), which we refer to as LoRA-P (P for preconditioning). We will show that both LoRA and LoRA-P benefit significantly from Nyström initialization. Due to space limitation, we summarize NoRA and NoRA+ in Algs. 1 and 2, respectively in the appendix, with additional explanations in Apdx. A.3.
@@
Deployment efficiency. NoRA offers practical advantages over other initialization methods such as PiSSA and OLoRA. It not only bypasses the computationally expensive SVD or QR decomposition, but also avoids the need to modify to the pretrained weights. NoRA is thus an off-the-shelf solution to enhance LoRA without altering existing pipelines. We expand on this in Apdx. A.3.
@@ A.1 More on related work
LoRA and parameter-efficient finetuning. LoRA (Hu et al., 2022) is a notable example of parameter-efficient finetuning (PEFT) (Dong et al., 2024) approaches. The goal of PEFT is to reduce the resource requirement for finetuning LLMs on downstream tasks. Other commonly adopted PEFT methods include, e.g., adapters (Houlsby et al., 2019), zeroth-order optimizers (Malladi et al., 2023; Zhang et al., 2024a, b), and prefix tuning (Li and Liang, 2021). There are also various efforts to further enhance LoRA via adaptivity (Zhang et al., 2023), chaining (Lialin et al., 2024; Xia et al., 2024), regularization (Li et al., 2024a; Li and Giannakis, 2023), low-bit training (Dettmers et al., 2023; Li et al., 2024b), modifications for long-sequences (Chen et al., 2024), weight decomposition (Liu et al., 2024), and combining with sparsity (Nikdan et al., 2024). Additionally, there are several approaches aiming at further reducing the number of trainable parameters in LoRA; examples include (Kopiczko et al., 2024; Lingam et al., 2024; Gao et al., 2024; Zhu et al., 2024; Hao et al., 2024; Bałazy et al., 2024). While originally designed for finetuning LLMs, LoRA also finds its applications in other domains, such as image generation (Gu et al., 2023) and continual learning (Smith et al., 2023).
@@ A.2 LoRA for linear models as asymmetric matrix factorization
A.3 More on NoRA and NoRA+

### v2

@@ 1 Introduction
ScaledGD (Tong et al., 2021) local ${\cal O}(\log(1/\epsilon))$
@@
OP modified GD* (Xiong et al., 2024) small ${\cal O}(\kappa^{2}\log(1/\epsilon))$
@@ 1.1 Related works
We only streamline results on the convergence of matrix factorization under the broad umbrella of optimization under forth order growth, where the objective is to minimize a forth-order polynomial with first order approaches. Other closely related topics, such as LoRA variants, can be found in Apdx. A.1.
Optimization under forth-order growth. Matrix factorization problems considered in this work are classical examples of forth-order growth functions. It involves a complex landscape characterized by nonconvexity, nonsmoothness, and the absence of PL condition. Similar to other works listed in Tab. 1, the goal of this work is to recap this classical problem and to unveil intriguing behaviors from an optimization perspective. Recent works have examined the convergence of several algorithms, such as GD, AltGD, and ScaledGD (Du et al., 2018; Ye and Du, 2021; Ward and Kolda, 2023; Jia et al., 2023; Jiang et al., 2023b) in exact- and over- parametrized settings. Most of them admit linear convergence with different dependences on the condition number of the factorized matrix $\mathbf{A}$. A concurrent work (xu2024provable) studies Nyström initialization for Nesterov’s accelerated method on the asymmetric and over-parametrized setting, giving an improved condition number dependence over GD. GD for matrix square root problems is studied in (Jain et al., 2017). Another closely related setting within forth-order growth is matrix sensing; see e.g., (Stöger and Soltanolkotabi, 2021; Zhang et al., 2021; Jin et al., 2023; Xiong et al., 2024; Xu et al., 2023). Linear rates are obtained for problems with exact- and over- parametrization, despite some of them demand early stopping. Similar to matrix factorization, not too much is known for under-parametrization. There are other approaches to tackle general forth-order growth optimization. For example, relative smoothness is considered in (Lu et al., 2018); adaptive step sizes induced by fine-grained geometry are studied in (Davis et al., 2024). The work of (Dragomir and Nesterov, 2023) also copes with such problems but requires convexity of the objective.
@@ 2.1 Preliminaries
Within this section, we assume that $\mathbf{A}\in\mathbb{R}^{m\times m}$ is positive semidefinite (PSD), otherwise one can employ the asymmetric formulation as in later sections. Problem (1) also closely links with matrix sensing, particularly under a sufficient number of Gaussian measurements (Xiong et al., 2024). From an optimization perspective, problem (1) is nonconvex and has no global Lipschitz gradient (Tu et al., 2016; Chi et al., 2019). These undesirable properties pose challenges for analyzing its convergence.
@@ 2.2 Nyström initialization
where $\mathbf{\Omega}\in\mathbb{R}^{m\times r}$ is a Gaussian random matrix. Note that this does not exclude other valid choices of $\mathbf{X}_{0}$, as long as conditions i) and ii) are satisfied. From this initialization, it is not difficult to see that condition i) is satisfied already. Our next lemma shows that the condition ii) holds w.h.p.
@@
The detailed expression for this “high probability” in Lemma 1 can be found in Apdx. B.1.1. Note that there will be a “w.h.p.” over the initialization in most of our results. This refers to that $\text{rank}(\mathbf{X}_{0})=r$ is needed for exact- and under-parametrized settings, and $\text{rank}(\mathbf{X}_{0})=r_{A}$ is needed when over-parametrized.
@@ 3.2 Nyström initialization in the exact-parametrized setting
With Nyström initialization (5) and different choices of step sizes, modified ScaledGD in (6) has a similar behavior as Theorem 1, that is, a two-phase behavior w.h.p. over the initialization:
@@ 4 NoRA: Nyström low rank adapters
where $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$ is the pretrained weight (of a particular layer), and $\mathbf{X}\in\mathbb{R}^{m\times r}$ and $\mathbf{Y}\in\mathbb{R}^{n\times r}$ with $r\ll\min\{m,n\}$. A more detailed recap of LoRA can be found in Apdx. A.1. Directional alignment can be achieved if singular vectors for $\Delta\mathbf{W}$ are leveraged to initialize $\mathbf{X}_{0}$ and $\mathbf{Y}_{0}$. While $\Delta\mathbf{W}$ is unavailable a priori, empirical wisdom suggests that there exists a set of well-performed adapters that lie in the column span of the pretrained weight matrix (Lingam et al., 2024), i.e., $\text{ColSpan}(\Delta\mathbf{W})\subseteq\text{ColSpan}(\mathbf{W}_{0})$. In other words, $\mathbf{W}_{0}$ can be adopted as a suitable replacement of $\Delta\mathbf{W}$ for directional alignment.
@@
We note that ScaledGD has already been applied for LoRA training in (Zhang and Pilanci, 2024), which we refer to as LoRA-P (P for preconditioning). We will show that both LoRA and LoRA-P benefit significantly from Nyström initialization. Due to space limitation, we summarize NoRA and NoRA+ in Algs. 1 and 2, respectively in the appendix, with additional explanations in Apdx. A.4.
@@
Deployment efficiency. NoRA offers practical advantages over other initialization methods such as PiSSA and OLoRA. It not only bypasses the computationally expensive SVD or QR decomposition, but also avoids the need to modify to the pretrained weights. NoRA is thus an off-the-shelf solution to enhance LoRA without altering existing pipelines. We expand on this in Apdx. A.4.
@@ A.1 More on related work
LoRA and parameter-efficient finetuning. LoRA (Hu et al., 2022) is a notable example of parameter-efficient finetuning (PEFT) approaches. The goal of PEFT is to reduce the resource requirement for finetuning LLMs on downstream tasks. Other commonly adopted PEFT methods include, e.g., adapters (Houlsby et al., 2019), zeroth-order optimizers (Malladi et al., 2023; Zhang et al., 2024a, b), and prefix tuning (Li and Liang, 2021). There are also various efforts to further enhance LoRA via adaptivity (Zhang et al., 2023), chaining (Lialin et al., 2024; Xia et al., 2024), regularization (Li et al., 2024a; Li and Giannakis, 2023), low-bit training (Dettmers et al., 2023; Li et al., 2024b), modifications for long-sequences (Chen et al., 2024), weight decomposition (Liu et al., 2024), and combining with sparsity (Nikdan et al., 2024). Additionally, there are several approaches aiming at further reducing the number of trainable parameters in LoRA; examples include (Kopiczko et al., 2024; Lingam et al., 2024; Gao et al., 2024; Zhu et al., 2024; Hao et al., 2024; Bałazy et al., 2024). While originally designed for finetuning LLMs, LoRA also finds its applications in other domains, such as image generation (Gu et al., 2023) and continual learning (Smith et al., 2023).
@@
Convergence of quasi-Newton methods. ScaledGD is sometimes regarded as a quasi-Newton method (Tong et al., 2021) for our nonconvex objective (1). Since much of the existing work on this topic focuses on the smooth and strongly convex case, we briefly review these approaches to highlight the significance of our results in achieving quadratic convergence for nonconvex and local-smooth problems. In the smooth and strongly convex regime, the primary advantage of quasi-Newton methods lies in their asymptotic ability to achieve super-linear convergence as $t\rightarrow\infty$. Non-asymptotic analyses demonstrating super-linear local rates have only been established recently; see, e.g., (rodomanov2021; jin2023non; ye2023towards; jiang2023online).
@@ A.2 LoRA for linear models as asymmetric matrix factorization
A.3 Initialization without knowing $\mathbf{A}$
In certain scenarios, direct access to $\mathbf{A}$ in problems (1) and (4) may not be available. Nevertheless, the Nyström initialization remains applicable, as it can be derived from the gradient, which is the minimum requirement for gradient-based methods. We take exact-parametrization as an illustrative example. For Nyström initialization (3) for the symmetric problem (1), we can simply calculate the gradient at $\mathbf{\Omega}$, i.e., $\mathbf{G}_{0}=(\mathbf{\Omega}\mathbf{\Omega}^{\top}-\mathbf{A})\mathbf{\Omega}$ and set initialization as $\mathbf{X}_{0}=-\mathbf{G}_{0}+\mathbf{\Omega}\mathbf{\Omega}^{\top}\mathbf{\Omega}$. For the asymmetric problem (4), our Nyström initialization in (5), can be obtained via the negative gradient at $\mathbf{X}=0$ and $\mathbf{Y}=\bf{\Omega}$.
A.4 More on NoRA and NoRA+